It's railed. This is what halted locking progess on Monday night, as this channel is used for the offloadMCF script, which slowly feeds back a CARM signal to the ETMs to prevent the VCO from saturating.
Attached is a 5 day trend, which shows that the channel went dead a few days ago. All the channels shown are being collected from the same ICS110B (I think), but only some are dead. It looks like they went dead around the time of the "All computers down" from Sunday.
Attached are the channels being recorded from the ICS110B in 1Y2 (the IOO rack). Channels 12, 13, 16, 17, 22, 24, 25 appear to have gone dead after the computer problems on Sunday.
This has been fixed by one of the two most powerful & useful IFO debugging techniques: rebooting. I keyed the crate in 1Y2.